# SceneTracker: Long-term Scene Flow Estimation Network

### v2

@@ Abstract
Considering the complementarity of scene flow estimation in the spatial domain’s focusing capability and 3D object tracking in the temporal domain’s coherence, this study aims to address a comprehensive new task that can simultaneously capture fine-grained and long-term 3D motion in an online manner: long-term scene flow estimation (LSFE). We introduce SceneTracker, a novel learning-based LSFE network that adopts an iterative approach to approximate the optimal trajectory. Besides, it dynamically indexes and constructs appearance and depth correlation features simultaneously and employs the Transformer to explore and utilize long-range connections within and between trajectories. With detailed experiments, SceneTracker shows superior capabilities in handling 3D spatial occlusion and depth noise interference, highly tailored to the LSFE task’s needs. Furthermore, we build the first real-world evaluation dataset, LSFDriving, to boost the LSFE field. The code for SceneTracker is available at https://github.com/wwsource/SceneTracker.
@@ 1 Introduction
    To boost the development of the LSFE field, we build the first real-world evaluation dataset, LSFDriving.
@@ 2.1 Scene Flow Estimation
The existing methods generally rely on feature similarities as their underlying principle. FlowNet3D [4] operates directly on point clouds with PointNet++ [5]. CamLiFlow [6] refines the optical flow and scene flow from the two frames’ feature similarity tensor in a coarse-to-fine way with two branches. OpticalExp [7] uses optical expansion to recover the pixel’s motion in depth between frames. It can estimate the final depth value of the second frame by combining the first frame’s depth, indirectly estimating the scene flow per pixel. The existence of occlusion issues severely constrains the accuracy of SFE. To tackle the occlusion problem, RigidMask [8] and RAFT-3D [9] focus on introducing rigid motion priors explicitly and implicitly, respectively.
@@ 5.5 Results on WWDriving
This subsection mainly analyzes the performance of our method on real-world data. TABLE III shows the evaluation results of our SceneTracker under different inference modes on the LSFDriving dataset after the Odyssey training process. We evaluate the three types of points in LSFDriving separately. We test the “Single” and “All” inference modes. In the “Single” inference mode, the network tracks only one query point at a time, while in the “All” inference mode, it tracks all query points simultaneously. Since each sample in LSFDriving only requires tracking up to five points, even tracking all query points simultaneously (“All” inference mode) does not provide the network with rich cross-space context. Therefore, we make the network incorporate additional auxiliary points, such as local points within the vicinity of the target point and global points uniformly distributed across the entire image, to track them together.
It can be seen from Table III that adding auxiliary points can significantly enhance the network’s performance, irrespective of the inference mode. Simultaneously, under the condition of adding auxiliary points, the “All” inference mode exhibits superior performance across all 3D dataset metrics. Furthermore, we observe that errors in pedestrian samples are consistently higher than those in background and vehicle samples. This fully showcases the complexity and challenges posed by non-rigid motions.
@@
5.6 Ablation Studies
Figure 5: Dataset evaluation metrics versus the number of FIM’s updates at inference time.
@@ 5.6 Ablation Studies
Number of FIM’s updates We train the model with 4 FIM’s updates and evaluate with $n=4$. The same inference update number gives the best results as evaluated in Fig. 5.
@@ 6 Conclusion
In summary, we propose a new task, i.e., , long-term scene flow estimation, to capture the fine-grained and long-term 3D motion better. To effectively tackle this issue, we introduce a novel approach named SceneTracker. With detailed experiments, SceneTracker shows superior capabilities in handling 3D spatial occlusion and depth noise interference, highly tailored to the LSFE task’s needs. To boost the development of the LSFE field, we build the first real-world evaluation dataset, LSFDriving.

### v3

@@ Abstract
Considering the complementarity of scene flow estimation in the spatial domain’s focusing capability and 3D object tracking in the temporal domain’s coherence, this study aims to address a comprehensive new task that can simultaneously capture fine-grained and long-term 3D motion in an online manner: long-term scene flow estimation (LSFE). We introduce SceneTracker, a novel learning-based LSFE network that adopts an iterative approach to approximate the optimal trajectory. Besides, it dynamically indexes and constructs appearance and depth correlation features simultaneously and employs the Transformer to explore and utilize long-range connections within and between trajectories. With detailed experiments, SceneTracker shows superior capabilities in handling 3D spatial occlusion and depth noise interference, highly tailored to the LSFE task’s needs. Finally, we build the first real-world evaluation dataset, LSFDriving, further substantiating SceneTracker’s commendable generalization capacity. The code and data for SceneTracker is available at https://github.com/wwsource/SceneTracker.
@@ 1 Introduction
    We build the first real-world evaluation dataset, LSFDriving, further substantiating SceneTracker’s commendable generalization capacity.
@@ 2.1 Scene Flow Estimation
The existing methods generally rely on feature similarities as their underlying principle. FlowNet3D [4] operates directly on point clouds with PointNet++ [5]. CamLiFlow [6] refines the optical flow and scene flow from the two frames’ feature similarity tensor in a coarse-to-fine way with two branches. OpticalExp [7] uses optical expansion to recover the pixel’s motion in depth between frames. It can estimate the final depth value of the second frame by combining the first frame’s depth, indirectly estimating the scene flow per pixel. The existence of occlusion issues severely constrains the accuracy of SFE. To tackle the occlusion problem, RAFT-3D [8] and RigidMask [9] focus on introducing rigid motion priors explicitly and implicitly, respectively.
@@ 5.5 Results on WWDriving
Figure 5: Examples of the proposed LSFDriving dataset.
This subsection mainly analyzes the performance of our method on real-world data. TABLE III shows the evaluation results of our SceneTracker under different inference modes on the LSFDriving dataset, whose examples are shown in Fig. 5, after the Odyssey training process. We evaluate the three types of points in LSFDriving separately. We test the “Single” and “All” inference modes. In the “Single” inference mode, the network tracks only one query point at a time, while in the “All” inference mode, it tracks all query points simultaneously. Since each sample in LSFDriving only requires tracking up to five points, even tracking all query points simultaneously (“All” inference mode) does not provide the network with rich cross-space context. Therefore, we make the network incorporate additional auxiliary points, such as local points within the vicinity of the target point and global points uniformly distributed across the entire image, to track them together.
It can be seen from Table III that adding auxiliary points can significantly enhance the network’s performance, irrespective of the inference mode. Simultaneously, the “All” inference mode exhibits superior performance across all 3D dataset metrics with auxiliary points. Furthermore, we observe that errors in pedestrian samples are consistently higher than those in background and vehicle samples. This fully showcases the complexity and challenges posed by non-rigid motions.
@@ 6 Conclusion
In summary, we propose a new task, i.e., , long-term scene flow estimation, to capture the fine-grained and long-term 3D motion better. To effectively tackle this issue, we introduce a novel approach named SceneTracker. With detailed experiments, SceneTracker shows superior capabilities in handling 3D spatial occlusion and depth noise interference, highly tailored to the LSFE task’s needs. Finally, we build the first real-world evaluation dataset, LSFDriving, further substantiating SceneTracker’s commendable generalization capacity.